four times the sum of a number and -52 amounts to -24
Found 2 solutions by algebrahouse.com, mananth:
Answer by algebrahouse.com(1659)

"four times the sum of a number and -52 amounts to -24"

x = the number

4x + (-52) = -24 {four times the sum of a number and -52 is -24}
4x - 52 = -24 {simplified the plus a negative}
4x = 28 {added 52 to both sides}
x = 7 {divided both sides by 4}

For more help from me, visit: www.algebrahouse.com



Answer by mananth(16949)

number =x
4times = 4(x+(-52))
4(x+(-52))=-24
4x-208=-24
add 208 to both sides
4x=208-24
4x=184
/4
x=46